# A fluid–peridynamic structure model of deformation and damage of microchannels

### v1

@@ B.1 Equations for the perturbations
First, substituting Eqs. (22) into the dimensionless peridynamic beam equation (16), we obtain and acknowledging that the base state (zero subscripts) solves the original problem, we obtain
@@ B.2 Transformation into a linear system
$Re\left[-i\omega\hat{Q}e^{i(kX-\omega T)}\right.\\ \left.+\frac{6}{5}ik\left(2Q_{0}\hat{Q}e^{i(kX-\omega T)}/H_{0}-Q_{0}^{2}\hat{H}e^{i(kX-\omega T)}/H^{2}_{0}\right)\right]\\ =-ikH_{0}\hat{P}e^{i(kX-\omega T)}-12\left(-2\frac{Q_{0}\hat{H}e^{i(kX-\omega T)}}{H^{3}_{0}}+\frac{\hat{Q}e^{i(kX-\omega T)}}{H^{2}_{0}}\right).$ (56)
By collecting like terms, we can rewrite the latter as By collecting like terms, the last equation can be rearranged into the final algebraic form:

### v2

@@ B.1 Equations for the perturbations
First, substituting Eqs. (22) into the dimensionless peridynamic beam equation (16), and acknowledging that the base state (zero subscripts) solves the original problem, we obtain
@@ B.2 Transformation into a linear system
$Re\left[-i\omega\hat{Q}e^{i(kX-\omega T)}\right.\\ \left.+\frac{6}{5}ik\left(\frac{2Q_{0}}{H_{0}}\hat{Q}e^{i(kX-\omega T)}-\frac{Q_{0}^{2}}{H^{2}_{0}}\hat{H}e^{i(kX-\omega T)}\right)\right]\\ =-ikH_{0}\hat{P}e^{i(kX-\omega T)}-12\left(-2\frac{Q_{0}\hat{H}e^{i(kX-\omega T)}}{H^{3}_{0}}+\frac{\hat{Q}e^{i(kX-\omega T)}}{H^{2}_{0}}\right).$ (56)
By collecting like terms, the last equation can be rearranged into the final algebraic form: